A rectangle has sides of 3x – 4 and 7x + 10. Find the expression that represents its perimeter.
[Perimeter = 2* width + 2*length]

Answer by checkley71(8403)

SO MULTIPLY EACH SIDE BY 2 & ADD:
2(3X-4)+2(7X+10)=6X-8+14X+20=20X+12=4(5X+3) IS THE PERIMETER.